find two consecutive even integers such that the lesser added to three times the greater is 46
Answer by checkley77(12844)

Let x & x+2 be the 2 even integers.
x+3(x+2)=46
x+3x+6=46
4x=46-6
4x=40
x=40/4
x=10 ans. for the smaller integer.
10+2=12 the larger integer.
Proof:
10+3*12=46
10+36=46
46=46